A sealed flask of capacity 500cm3 contains gas at a pressure of 1 atmosphere. If the temperature of the gas is raised from 10℃ to 35℃,without any change in volume, calculate the new pressure.
1
Expert's answer
2020-05-26T12:49:44-0400

We can use the combined gas law (CGL) obtained from the ideal gas law. For two different states 1 and 2 of gas in an isolated system, the CGL states:

"\frac{p_1V_1}{T_1}=\frac{p_2V_2}{T_2}."

The pressure can be in Pa, atm, or bar, the temperature must be in K. Since the volume did not change, or "V_1=V_2", we can re-write the equation:

"\frac{p_1}{T_1}=\frac{p_2}{T_2},
\\\space\\
p_2=p_1\frac{T_2}{T_1}
\\\space\\
p_2=1\frac{273+35}{273+10}=1.09\text{ atm}."